# revtex4-2Repair the float

### v1

@@ Abstract
Lying at the interface between Network Science and Machine Learning, node embedding algorithms take a graph as input and encode its structure onto output vectors that represent nodes in an abstract geometric space, enabling various vector-based downstream tasks such as network modelling, visualization, data compression, node classification, link prediction, and community detection. Two apparently unrelated limitations affect these algorithms. On one hand, it is not clear what the basic operation defining vector spaces, i.e. the vector sum, corresponds to in terms of the original nodes in the network. On the other hand, while the same input network can be represented at multiple levels of resolution by coarse-graining the constituent nodes into arbitrary block-nodes, the relationship between node embeddings obtained at different hierarchical levels is not understood. Here, building on recent results in network renormalization theory, we address these two limitations at once and define a multiscale node embedding method that, upon arbitrary coarse-grainings, ensures statistical consistency of the embedding vector of a block-node with the sum of the embedding vectors of its constituent nodes. We illustrate the power of this approach on two economic networks that can be naturally represented at multiple resolution levels: namely, the network of international trade between (sets of) countries and the network of input-output flows among (sets of) industries in the Netherlands. We confirm the statistical consistency between networks retrieved from coarse-grained node vectors and networks retrieved from sums of fine-grained node vectors, a result that cannot be achieved by alternative methods. Several key network properties, including a large number of triangles, are successfully replicated already from embeddings of very low dimensionality, allowing for the generation of faithful replicas of the original networks at arbitrary resolution levels.
@@ I Introduction
Complex networks capture a variety of socially relevant processes, from economic activities to interactions among brain regions [1, 2, 3]. Indeed, every dyadic interaction can be described by properly defining which are the actors (nodes) and the type of connections among them (edges). By defining the nodes as the sectors and the edges as the transactions among them, the Input-Output network (ION) [4] is recovered. Again, by defining the nodes as the states and the edges as the trade among them, the World Trade Web (WTW) [5] is obtained. On the other hand, if nodes were seen as the brain regions connected by electrical stimuli, the brain network is represented. In particular, this flexibility allows for arbitrary definitions of the nodes even when looking at the same phenomenon that is generating the data (generative process). For example, one could have access to the network involving the most detailed sector classification (National Industry), whereas another one only at the community level (Industry) - see the upper part of Figure 1. The same reasoning applies naturally in other contexts (neuroscience, social sciences, $\cdots$), as community structures help in simplifying the heterogeneity of the graph [6, 7]. Therefore, seen at a coarser resolution, the graph represents the interactions among block-nodes, and it would be uniquely recovered after the specification of the partitions. This scheme could be iterated at wish to produce a multi-scale unfolding of the original graph with nested partitions: pictorially, the base of a pyramid is the observed network, whereas the coarser levels are the cross-sections of the pyramid. Lastly, it is worth noticing that the properties of each lower-resolution graph change with levels. For instance, the firm graph is less dense than the sector one since there will be fewer nodes to redistribute links to.
@@
Here, assuming to know the nodes, we aim at modeling their interactions by assigning a probability for every pair of nodes (or edge or link) [8]: the higher the probability, the more likely is that edge to exist in a sampled graph. Ultimately, the measurements over the observed graph should coincide with the average over the sampled networks. This exercise is called, in general, network modeling but also network reconstruction [9] or (binary) edge classification [10] in the machine-learning literature.
To tackle this problem, many machine-learning models use node embeddings [11, 12, 2]. As highlighted by the arrows in Figure 1, one may assign a set of coordinates for every node and, then, extract the probability for the network’s edges. The optimal node embeddings are the vector of parameters that optimize a functional involving the observed graph and the model, for e.g. the likelihood. Therefore, these vectors are interpreted as informative features about the nodes that can be deployed to different tasks, such as community detection or node classification[13].
@@
Two hallmarks of real-world graphs are 1) low density (sparsity) and 2) high triangle density, where there are many triangles incident to low-degree vertices [14]. It was recently proven that linear node embedding methods, such as node2vec [11], are not capable of reproducing the triangle density [14]. To overcome this drawback, LogisticPCA [15] optimized the vectors for a non-linear logistic function. Furthermore, Chanpuriya et colleagues proposed the “symmetric” LogisticPCA (LPCA) [16] to deal with undirected networks (see subsection III.1).
@@
As mentioned, a phenomenon can be studied at different scales. By combining nodes into communities we go from a microscopic to a coarse grained scale. Nevertheless, most of the models, e.g. LPCA, regard a network only at one scale, providing the optimal embeddings for that level. If nodes were merged into communities, the block-vectors have to be recomputed. In other words, the two sets of vectors are completely unrelated, as if the models see the two networks as realizations of two distinct processes - even though it is not the case. there is no prescription to use the node embeddings from a lower level to create an embedding for a community. For this reason, we would refer to this class of models as “single-scale” models (SSM).
@@
To tackle the renormalization problem on networks, several methods have been proposed in the literature, but they all rely on strong assumptions that limit their use cases [17]. More concretely, the most promising one [3] assumes that the nodes are embedded in a hyperbolic plane, all the coarser networks are scale-free and the block-nodes contain $r=2$ micro-nodes. The latter restriction doesn’t allow to aggregate the nodes with the “most natural” way induced by the studied phenomenon, e.g. geographical distances for the WTW or the sector (industry) for the ION. To overcome this limitation, the multi-scale model [17] was proposed, which is generalizable at higher levels and allows for any arbitrary partition of the microscopic nodes. Indeed, the block parameters are uniquely obtained by summing the vectors of the nodes belonging to a community (renormalization rule) - this mimics the unique identification of the coarser network starting from the microscopic one.
@@
In this work, we enhance the scalar multi-scale model with node embeddings (MSM). The renormalization rule for vectors is shown in Figure 1: each community embedding is the summation of its lower-level ones. Therefore, the MSM provides an interpretation of the sum of node embeddings, which is rarely addressed in the literature [^1] The successes of Natural Language Processing are also due to the effective representation of a phrase obtained by summing the embedding for each word of it [18]. This is hardly replicated in a graph setting as it is intrinsically more complex than a language. [19]. Due to the renormalization rule, the MSM has the additional benefit of having to be fitted only at the ground level, further implying a lower computational complexity with respect to a single-scale model to be tailored at every level.
@@
The rest of the paper is organized as follows. In section III, we introduce the LogisticPCA and the Multi-Scale Model (MSM) alongside its renormalization rule. In subsection IV.1, we describe the ING Input-Output Network and the World Trade Web datasets. Furthermore, we present the coarse-graining procedure to obtain the higher-scale network. In section VI, we show the multi-scale results of the two models and discuss the implications of the theoretical results over either network- and machine-learning scores. Finally, in the Appendices, we store all the technical details supporting the results in the main text.
Figure 1: Visualization of the research question: how node embeddings relate across scales? Here, it is shown the embedding extraction scheme for two levels: on the left the microscopic level that uniquely generates the coarser one on the right after aggregation. However, for SSM, the vectors fitted at the microscopic level can’t be used to obtain the ones computed at level 1 (red question marks), whereas its the case for our MSM. The latter succes is possible thanks to the vector sum which is highlighted by using the parallelogram law on the right, e.g. pale red vectors are aligned in order to get the solid red vector at an higher scale.
@@ II Graph Renormalization
In this section, we will introduce the mathematical description of a graph and its coarse-graining procedure. We use the subscript $\ell=0$ to identify the base graph whereas $\ell>0$ for its aggregated versions. In addition, we refer to a generic quantity at level $\ell$ as $\ell$-quantity, e.g. the 2-vectors and 2-nodes are, respectively, the node embeddings and block-nodes at level $2$.
Consider a binary undirected graph with $N_{0}$ “microscopic” nodes $\mathcal{V}_{0}$ (labelled as $i_{0}\in[1,N_{0}]$) and their connections (called edges or links), i.e.
@@
where $a^{(0)}_{i_{0}j_{0}}=1$ if there was an edge among $i_{0},j_{0}$ and $a^{(0)}_{i_{0}j_{0}}=1$ otherwise. This system could be represented by an $N_{0}\times N_{0}$ adjacency matrix $\mathbf{A}^{(0)}$ which has to be symmetric, i.e. $a^{(0)}_{i_{0}j_{0}}=a^{(0)}_{j_{0}i_{0}}$, since the graph is undirected. We don’t account for multiple edges, whereas we do for self-loops in the diagonal of $\mathbf{A}^{(0)}$, i.e. $a^{(0)}_{i_{0}i_{0}}=0,1$. Unless explicitely specified, we would use the notation $\xi_{ij}:=\xi^{(\ell)}_{i_{\ell}j_{\ell}}$ with $\xi$ as a chosen quantity when the results are valid for any level $\ell\geq 0$, e.g. if $\xi:=\mathbf{A}^{(\ell)}\textnormal{ then }a_{ij}:=a^{(\ell)}_{i_{\ell}j_{\ell}}$.
@@ II.1 Coarse-Graining
In order find the coarse-grained version of the graph, we define non-overlapping arbitrary partitions $\Omega_{0}$ of the microscopic nodes into block-nodes
@@
Secondly, we assume that two communities are connected if there was at least one edge between their internal nodes, i.e.
@@
where the operation used is the logical OR over the edges connecting two communities. That is, only the zeros are preserved whereas the possible multiple-edges among the lumped nodes are projected to one. Moreover, note that $\Omega_{0}$ is only surjective but not injective, since multiple nodes $i_{0}$ are merged into $i_{1}$, and that we allow for self-loops as we didn’t require $i_{1}\neq j_{1}$. At a higher level, the self-loop proxies there is at least one connection inside the community either because links among nodes or self-loops edges. In this way, we built the $N_{1}\times N_{1}$ adjacency matrix $\mathbf{A}^{(1)}$ which is binary and symmetric as $\mathbf{A}^{(0)}$. To ease the recalling, we will refer to $A^{(0)}$ as 0-graph, $A^{(1)}$ as 1-graph, the $N_{0}$ microscopic nodes as $0-nodes$ and the block-nodes as $1-nodes$.
The lumping procedure could be iterated $\ell+1(\geq 1)$ times by specifying the partition $\Omega_{\ell}$ of $\ell-nodes$ into $N_{\ell+1}$ block-nodes
@@
However, since the partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ are not overlapping, one can define their composition $\Omega_{0\to\ell}$ as
$\Omega_{0\to\ell}:=\Omega_{\ell}\circ\,\cdots\,\circ\Omega_{0}.$ (2)
This provides a direct mapping of $0-nodes$ into block-nodes $i_{\ell+1}=\Omega_{0\to\ell}(i_{0})$. Therefore, the $(\ell+1)-graph\textnormal{, }\mathbf{A}^{(\ell+1)}$ is recovered both applying iteratively the Equation 1 or directly via Equation 2
$a^{(\ell+1)}_{i_{\ell+1}j_{\ell+1}}$ $=1-\prod_{i_{\ell}\in\Omega_{\ell}^{-1}(i_{\ell+1}),j_{\ell}\in\Omega_{\ell}^{-1}(j_{\ell+1})}(1-a^{(\ell)}_{i_{\ell}j_{\ell}})$ (3)
@@
The nested $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ can be uniquely parametrized in terms of a dendrogram as shown in [17]. By following an “horizontal” cut of the dendrogram, one obtains the community partitions at the same scale, e.g. for the WTW by merging states nearer than a fixed geographical distance. In contrast, by cutting the dendrogram at different heights, one recovers the “multiscale” clusters, e.g. a state interacting with a continent.
In many cases, the problem setting suggests a partitioning of the nodes into communities without an explicit distance matrix, e.g. the classification of the Industries based on the NAICS-codes digits. They characterize each Industry by a number at $6$ digits, e.g. Full-Service Restaurants (722511), but also their clusters by progressively removing the last number up to $2$ digits, e.g. Accommodation and Food Services (72). More concretely, the industries sharing the first $n\in[2,5]$ digits are merged into a block-node at the $\ell:=6-n\in[1,4]$ level (cfr. Hamming Distance). The dendrogram can be built by setting the height where the blocks split into sub-cluster as $\ell$.
Lastly, if the problem doesn’t come with any “natural” partition, it would be possible to merge the 0-nodes randomly by creating a dendrogram from a random distance matrix.
III Models
In the previous sections, we highlighted that one phenomenon could be studied at different resolutions. In particular, since every adjacency matrix $\mathbf{A}^{(\ell)}$ is binary and undirected, every edge $(i_{\ell},j_{\ell})$ could be seen as a Bernoulli random variable. Since the theory is valid for every scale, we will remove the dependence on the level $\ell$ on each quantity, e.g. $a_{ij}:=a^{(\ell)}_{i_{\ell}j_{\ell}}$ (see section II). Hence, for every resolution the adjacency entries are

### v2

# Multi-Scale Node Embeddings for Graph Modeling and Generation

@@ Abstract
Lying at the interface between Network Science and Machine Learning, node embedding algorithms take a graph as input and encode its structure onto output vectors that represent nodes in an abstract geometric space, enabling various vector-based downstream tasks such as network modelling, data compression, and community detection. Two apparently unrelated limitations affect these algorithms. On one hand, it is not clear what the basic operation defining vector spaces, i.e. the vector sum, corresponds to in terms of the original nodes in the network. On the other hand, while the same input network can be represented at multiple levels of resolution by coarse-graining the constituent nodes into arbitrary block-nodes, the relationship between node embeddings obtained at different hierarchical levels is not understood. Here, building on recent results in network renormalization theory, we address these two limitations at once and define a multiscale node embedding method that, upon arbitrary coarse-grainings, ensures statistical consistency of the embedding vector of a block-node with the sum of the embedding vectors of its constituent nodes. We illustrate the power of this approach on two economic networks that can be naturally represented at multiple resolution levels: namely, international trade between (sets of) countries and input-output flows among (sets of) industries in the Netherlands. We confirm the statistical consistency between networks retrieved from coarse-grained node vectors and networks retrieved from sums of fine-grained node vectors, a result that cannot be achieved by alternative methods. Several key network properties, including a large number of triangles, are successfully replicated already from embeddings of very low dimensionality, allowing for the generation of faithful replicas of the original networks at arbitrary resolution levels.
Complex Systems — Statistics — Machine Learning
@@ I Introduction
Complex networks provide a powerful framework for modeling a variety of socially relevant processes, from economic activities to interactions among brain regions [1, 2, 3]. Essentially, any (dyadic) interaction can be represented as a network, by properly defining the entities as nodes and their interactions as edges. For example, the transactions between the economic sectors are represented by the Input-Output network (ION) [4], which interprets the former as edges and the latter as nodes. Similarly, the World Trade Web (WTW) [5], which tracks the trade between countries, assumes the countries as nodes and the trade flows as edges.
@@
The downside of this flexibility is that, even when looking at the same generative process, there is no unique definition of the nodes involved in the interaction. By recalling the ION, one may have access to the transactions involving the most detailed sector classification (National Industry), whereas another practitioner to a coarser version of it (Industry) - see Figure 1. As expected, the networks at these two resolutions are likely to have different properties. However, they should be described by a model consistently connecting the two scales, since the original Industry level was obtained by aggregating the National Industry one. In general, at a coarser resolution, the network represents the interactions among block-nodes, and it is uniquely recovered after specifying the partition $\Omega$ of the microscopic nodes into blocks. This scheme could be iterated at wish to produce a multi-scale representation of the original graph with nested partitions: pictorially, this can be thought as a pyramid where the base is the observed network and higher levels are the coarser scale with a lower and lower number of node up to the unique vertex. Lastly, it is worth noticing that the properties of each low-resolution graph change with levels. For instance, the firm graph is less dense than the sector one since there will be fewer nodes to redistribute links to. The same argument applies to other domains (neuroscience, social sciences, $\cdots$), as community structures are widely used to simplifying the heterogeneity of a graph [6, 7]. Therefore, the methodologies exposed in this essay can be applied beyond the examples, we are going to show in the Data and Results sections.
@@
In general, the nodes of a network are known (firms in a country, regions in the brain, people in a social network, $\dots$), whereas their observed interactions are a single realization of a hidden process. Therefore, by treating the nodes as fixed entities, we aim at modeling their interactions by assigning a probability for every pair of nodes (or edge or link) [8]: the higher the probability, the more likely is the edge to exist in a sampled graph. Ultimately, the measurements over the observed graph should coincide with the average over the sampled networks. This exercise is called, in general, network modeling, but also in the machine-learning literature network reconstruction [9] or (binary) edge classification [10].
The machine-learning models tackle this problem by exploiting the node embeddings [11, 12, 2] which are, in essence, the node parameters on which the probability function depends. Their values are obtained by maximizing the likelihood, or more in general by any other functional optimization. In particular, the general scheme is the one depicted in Figure 1. One starts by setting the initial conditions on node embeddings (“embedding map”). Then, the inverse map returns the connection probabilities, giving rise to the likelihood for that set of parameters. In turn, the latter produces the parameter updates towards the optimum. As said, this back-and-forth procedure ends when the likelihood is maximal, or the functional is at an optimum. The vectors are commonly regarded as informative “features” to be deployed in different tasks, such as link prediction or community detection[13, 14, 15].
@@
Two hallmarks of real-world graphs are 1) low density (sparsity) and 2) high triangle density - many triangles incident to low-degree vertices. Meet them both is challenging, as proven in [16]. In particular, every linear node embedding methods, such as node2vec [11], is not capable of reproducing the triangle density of a real network with a low embedding dimension -even with an a posteriori application of a non-linear activation function. Nevertheless, the problem can be solved by directly maximizing the likelihood of a non-linear logistic function - see LogisticPCA (LPCA) [14].
@@
As mentioned, a phenomenon may be studied at different scales depending on the resolution level one has access to (see Figure 1). Specifically, by combining nodes into communities, one goes from a microscopic to coarse grained scales. Nevertheless, the majority of the models, e.g. LPCA, regard a network only at one scale, providing the optimal embeddings for that level. If nodes were merged into communities, the block vectors have to be re-fitted, implying that these two sets of vectors, e.g. at level $\ell=\left\{0,1\right\}$, happened to be completely unrelated as if the models were describing two distinct processes - even though we know it was not the case. Moreover, if one wants to re-use the lower-level embeddings, there is no prescription to create an embedding for a community (see subsection A.2). For this reason, we are going to call this class of models as “single-scale” models (SSM).
@@
Several methods have been proposed in the literature, to solve the multi-scale-classification problem, but they all rely on strong assumptions that limit their use cases [17]. The most promising one [3] assumes that the nodes are embedded in a hyperbolic plane, all the coarser networks are scale-free and the block-nodes contain $r$ micro-nodes. The latter restriction doesn’t allow to aggregate the nodes with the “most natural” partition induced by the studied phenomenon, e.g. geographical distances for the WTW or the sector (industry) for the ION. To overcome these limitations, the multi-scale model was designed [17] to be generalizable at higher levels, whereas allowing for any arbitrary partition of the microscopic nodes. Specifically, for each block, its scalar parameters, or fitnesses, is uniquely obtained by summing the fitnesses of its internal nodes (renormalization rule). This operation can be seen as the counterpart, in the parameter space, of the aggregation procedure. Note that also the latter operation uniquely identifies a community starting from its microscopic members.
@@
In this work, we equip the scalar multi-scale model with node embeddings (MSM), for which the summation rule on vectors is shown in Figure 1. The benefits are two-folds. On the one hand, it provides an interpretation of the sum of node embeddings, which is rarely addressed in the literature [^1] The successes of the Natural Language Processing are also due to the effective representation of a phrase obtained by summing the embedding for each word[18]. This is hardly replicated in a graph setting, as it is more complex than a language [19]. [15]. The MSM, instead, assumes that the community vector is the sum of the embeddings of its members. Therefore, the node parameters are now used as vectors, that can be summed as in a proper vector space. The second benefit is that the MSM can be fitted even only at the microscopic level, since the coarser-block embeddings can be obtained under summation. In other words, there is no need to re-fit the model at higher scales. This further implies a lower-computational complexity with respect to a single-scale model that has to be refitted at every level.
@@
The rest of the paper is organized as follows. In section III, we introduce the LogisticPCA and the Multi-Scale Model (MSM) alongside its summation rule. In subsection IV.1, we describe the ING Input-Output Network and the World Trade Web datasets. Furthermore, we present the coarse-graining procedure to obtain the higher-scale network. In section V, we show the multi-scale results of the two models and discuss the implications of the theoretical results over either network- and machine-learning scores. Finally, in the Supplemental Material [20], we store all the technical details supporting the results in the main text.
Figure 1: The ground truth appears at multiple scales, depending on resolution of the dataset. Nevertheless, The macro-scale representation (shown in the plots on the right) can be uniquely obtained by coarse-graining the network at the microscopic scale. But how node embeddings are connected across different scales? On the left, one finds the learning procedure of the microscopic node embeddings, whereas on the right, the macroscopic counterparts. In single-scale models, the micro-vectors cannot be directly used to calculate the macro-vectors (indicated by the red question marks). In contrast, the multi-scale model overcomes this limitation, as the macro-vectors are the sum of micro-embeddings (see the parallelogram law on the right).
@@ II Graph Renormalization
In this section, we introduce the mathematical framework for representing a graph and its coarse-graining procedure. The subscript $\ell=0$ identifies the microscopic graph, while $\ell>0$ refers to its aggregated versions. We refer to any quantity defined at level $\ell$ as an $\ell$-quantity. For instance, the 2-vectors and 2-nodes denote the node embeddings and block-nodes at level $2$, respectively.
Let us consider a binary, undirected graph with $N_{0}$ “microscopic” nodes $\mathcal{V}_{0}$ (indexed by $i_{0}\in[1,N_{0}]$) and the connections among them (called edges or links), defined as
@@
where $a^{(0)}_{i_{0}j_{0}}=1$ indicates the present of an edge between the nodes $i_{0},j_{0}$ and $a^{(0)}_{i_{0}j_{0}}=1$ otherwise. This system is represented by an $N_{0}\times N_{0}$ symmetric matrix $\mathbf{A}^{(0)}$, such that $a^{(0)}_{i_{0}j_{0}}=a^{(0)}_{j_{0}i_{0}}$, due to the undirected nature of the graph. We do not allow for multiple edges, but self-loops are permitted, and they are stored in the diagonal entries of $\mathbf{A}^{(0)}$, i.e. $a^{(0)}_{i_{0}i_{0}}=0,1$.
@@ II.1 Coarse-Graining
To construct a coarse-grained version of the graph, we define a non-overlapping arbitrary partition $\Omega_{0}$ of the microscopic nodes into block-nodes (communities)
@@
Secondly, two blocks are connected if there was at least one edge between any pair of their constituent nodes. Formally,
@@
where $\Omega_{0}^{-1}(k)$ is the set of microscopic nodes grouped into block-node $k$. This operation effectively performs a logical OR across all edges connecting two communities. That is, only the zeros are preserved, whereas the multiple microscopic edges are seen as one interaction among the two groups.
Note that $\Omega_{0}$ is a surjective function: multiple nodes $i_{0}$ belong to exactly one block $i_{1}$. Self-loops are also allowed at the coarse-grained level ($i_{1}=j_{1}$), revealing that there was at least one intra-cluster connection -either an interaction between two different nodes or a self-loop at the microscopic level. The process returns a new $N_{1}\times N_{1}$ adjacency matrix $\mathbf{A}^{(1)}$ which remains binary and symmetric, as $\mathbf{A}^{(0)}$. For reference, we denote $A^{(0)}$ as 0-graph, $A^{(1)}$ as 1-graph, and refer to their nodes as 1nodes1-nodes and 0nodes0-nodes, respectively.
II.1.1 Recursive Coarse-Graining
This coarsening scheme may be recursively applied. At each level $\ell+1(\geq 1)$, we define a new partition $\Omega_{\ell}$ that merge $\ell$-nodes into $N_{\ell+1}$ blocks, namely
@@ II.1.1 Recursive Coarse-Graining
Given a sequence of non-overlapping partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$, one can define their composition $\Omega_{0\to\ell}$ as
$\Omega_{0\to\ell}:=\Omega_{\ell}\circ\,\cdots\,\circ\Omega_{0},$ (2)
which provides a direct mapping from micro 0-nodes to block-nodes $i_{\ell+1}=\Omega_{0\to\ell}(i_{0})$. Therefore, the $(\ell+1)$-graph $\mathbf{A}^{(\ell+1)}$ may be recovered either iteratively by means of Equation 1, or directly via
[…]
@@
The collection of nested partitions $\left\{\Omega_{\ell}\right\}_{\ell\geq 0}$ can be uniquely represent with a dendrogram as shown in [17]. A “horizontal” cut through the dendrogram yields the community partitions at the same scale, whereas, by cutting the dendrogram at different heights, it yields the “multiscale” clusters. More concretely, the two schemes for the World Trade Web (WTW) consist of merging states nearer than a fixed geographical distance (horizontal clustering) or a state with a continent (multiscale clustering).
In many real-world applications, node partitions are naturally suggested by the problem context, even when an explicit distance metric is missing. An example is the NAICS classification [21] of sectors, where each national industry (NI) is identified by a number at $6$ digits, e.g. Full-Service Restaurants (722511). By removing the last digit, one groups the $6$-digits NI into less-detailed descriptions up to $2$ digits, e.g. Accommodation and Food Services (72). Specifically, sectors sharing the first $n\in[2,6]$ digits are grouped into blocks at level $\ell:=6-n\in[0,3]$. Interestingly, this induces a Hamming distance between industries, defined as the number of differing digits. Still, this example highlighted that every partition is possible. In the absence of a natural partition, one can opt either for community detection algorithms Louvain [7, 22] or merge at random the 0-nodes using a random distance matrix among them.
III Methodology
In the previous sections, we emphasized that a single phenomenon can be analyzed at multiple levels of resolutions. More precisely, since every adjacency matrix $\mathbf{A}^{(\ell)}$ is binary and undirected, each edge $(i_{\ell},j_{\ell})$ can be modeled as a Bernoulli random variable.
To simplify the notation, we introduce the shorthand $\xi_{ij}:=\xi^{(\ell)}_{i_{\ell}j_{\ell}}$ to denote a generic quantity $\xi$ at level $\ell\geq 0$, in cases where results hold across all levels. For instance, if $\xi:=\mathbf{A}^{(\ell)}$, then $a_{ij}:=a^{(\ell)}_{i_{\ell}j_{\ell}}$. Hence, for every level $\ell$, the edge variable $a_{ij}$ distributes according to